how to solve using linear combination's 2x-y=8 and x+y=1


Answer by stanbon(48568)

how to solve using linear combination's
2x-y=8
x+y=1
-------------------------
Multiply thru the 2nd equation by 2 to get:
2x-y = 8
2x+2y = 2
------------------------------
Subtract the 1st from the 2nd Eq:
3y = -6
y = -2
------------------
Substitute into x+y = 1 to solve for "x":
x + -2 = 1
x = 3
---------------------------
Cheers,
Stan H.